# Differentially Private Knowledge Distillation via Synthetic Text Generation

### v2

@@ 1 Introduction
Large Language Models (LLMs) have accomplished amazing feats, from generating human-like text to improving human productivity through AI-powered assistants. However, recent work has demonstrated that practical privacy attacks are possible due to memorization of training data from LLMs Carlini et al. (2019, 2021). Differential Privacy (DP), a rigorous mathematical framework, is widely used to mitigate this type of information leakage Dwork (2006). To guarantee DP for machine learning models, DP is operationalized within the learning algorithm, known as DP-SGD Abadi et al. (2016). The standard training paradigm of LLMs is to first pre-train an LLM on large publicly available data, then fine-tune it on a specific downstream task using private data. Recent works in differentially private training have taken advantage of this paradigm by initializing an LLM using pre-trained weights and then applying DP-SGD in the fine-tuning process, which can significantly boost its performance Li et al. (2021); Yu et al. (2021); Ganesh et al. (2023), albeit with some reduction in model utility.
Due to the huge computational and memory demands of LLMs, it becomes necessary to consider model compression in this training paradigm. For example, model compression is required when deploying LLMs on client devices where memory resources are scarce, or real-time systems such as sentence completion in text editors where inference latency is crucial Xu et al. (2023). One popular technique for model compression that our work focuses on is task-specific Knowledge Distillation (KD) Hinton et al. (2015), which distills the knowledge of a larger teacher model onto a compressed student model. Although knowledge distillation in LLMs has been quite successful Jiao et al. (2019); Sanh et al. (2019); Sun et al. (2019), few works have explored compressing LLMs with DP Mireshghallah et al. (2022); Yu et al. (2023a, b). Since achieving DP or KD for an LLM already trade-offs utility loss, accomplishing both can result in substantial utility loss, which highlights the challenge of this problem.
@@
A naive solution to this problem would be to apply DP-SGD when training the teacher model, then apply DP-SGD again during the knowledge distillation between the teacher and student. Since the teacher model is a function of the entire private corpus, applying DP-SGD during the fine-tuning of the teacher is seemingly unavoidable. Despite the additive computational and memory costs due to an additional run of DP-SGD, this solution has also been shown to have worst utility compared to differentially private fine-tuning the student without knowledge distillation Mireshghallah et al. (2022).
@@ 2 Related Works
Knowledge Distillation and DP Compression. Work in knowledge distillation for LLMs typically trains a smaller student model to match its output logits with the output logits of a larger, teacher model. More advanced works align the hidden state vectors between the teacher and the student Jiao et al. (2019); Sanh et al. (2019); Sun et al. (2019).
@@
DP-SGD Abadi et al. (2016) is the most widely used technique to guarantee DP for LLMs. Specifically, a pre-trained LLM trained on large amounts of public text data scraped from the internet is further fine-tuned on a private corpus $D$ using DP-SGD McMahan et al. (2017); Yu et al. (2021); Li et al. (2021). An orthogonal approach investigated differentially private decoding of LLMs Flemings et al. (2024); Majmudar et al. (2022); Ginart et al. (2022), thereby avoiding the use of DP-SGD.
@@
Mireshghallah et al. (2022) first introduced DP model compression and gave a framework for DP knowledge distillation. However, their framework requires using DP-SGD twice: once for the teacher and once for the student. This procedure leads to suboptimal results for text classification tasks. Wang et al. (2019) also explored private model compression, but is limited by the number of queries to the teacher for knowledge distillation. Yu et al. (2023b) explores DP model compression during the pre-training stage; our work focuses solely on compression during fine-tuning. Another line of work similar to our approach is Data-Free Distillation Chen et al. (2019); Haroush et al. (2020) where a generative model synthesizes data to substitute training data. Since no private data is used, this results in no privacy loss, $\epsilon=0$. Our approach uses a private corpus to generate synthetic data in a differentially private manner.
@@
Another problem setup broadly related to DP knowledge distillation is private ensemble learning via PATE Papernot et al. (2016, 2018). An ensemble of teacher models is trained on disjoint private data, then produces noisy labels on unlabeled public data, which comes from a similar distribution as the private data, for the student to train on. Lyu and Chen (2020) used PATE to achieve differentially private knowledge distillation for mobile analytics. Our method does not require unlabeled public data.
@@
DP Synthetic Data for Downstream Task. Promising progress has been made in generating high-fidelity, DP synthetic text Bommasani et al. (2019); Mattern et al. (2022); Kurakin et al. (2023); Yue et al. (2022). These works have substantially reduced the utility gap between DP-SGD on real data and SGD on DP synthetic data for classification tasks.
@@
Yu et al. (2023a) investigated synthetic data generation from DP foundational models to train lightweight downstream models. However, our method differs from theirs by the following: (1) Their framework only considers knowledge transfer from the synthetic data. Our results demonstrate that other forms of knowledge distillation, in particular, the output distribution and hidden representations, are crucial to obtain strong performance for compressed models. (2) Their framework needs large foundational models to generate high-enough quality synthetic text data, which requires model parallelism on four V100 GPUs for fine-tuning even with LoRA Hu et al. (2021); Yu et al. (2021). And to the best of our knowledge, there is no open-source implementation of DP model parallel training. In contrast, our framework can obtain strong results without relying on large foundational models and thus does not need model parallelism.
@@ 3.1 Knowledge Distillation
Knowledge Distillation (KD) aims at transferring the knowledge of a teacher LLM $T$ to a smaller student LLM $S$. The standard approach is to train the student to mimic the output distribution of the teacher and the true labels. Let $f_{T}(x)$ and $f_{S}(x)$ be the logit outputs of the teacher and student networks evaluated on an input data $x$, respectively. $P_{T}$ and $P_{S}$ be the softmax outputs of the teacher and student networks, respectively, $Z_{T}$ and $Z_{S}$ be the logit outputs of the teacher and student networks, respectively, and $\mathbf{y}$ be the true labels. The student is trained to minimize the linear combination of the supervised learning loss and the distillation loss:
@@
where $\lambda$ is a weight hyperparameter, $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, $\mathcal{L}_{\text{KL}}$ is the Kullback-Leibler divergence, $\sigma$ is the softmax function, and the second term is the distillation loss which follows the softmax-temperature using a temperature parameter $t$ Hinton et al. (2015). At inference time, $t$ is set to 1 to recover the standard softmax. y is known as the hard label because of its one-hot encoding while $\sigma(f_{T}(x)/t)$ is the soft label.
@@ 3.2 Text Generation
Given some context vector $\mathbf{x}_{t}=x_{1},x_{2},...,x_{t}$, which is a string of tokens from some vocabulary $V$, i.e. $x_{i}\in V$ for all $i=1,...,t$, text generation involves sampling the next token $x_{t+1}$ using a generative language model $p$. More precisely, the output of a language model $p$ for a given context $\mathbf{x}_{t}$ is a likelihood function of all possible tokens $p(x_{t+1}=w|\mathbf{x}_{t})$, and choosing the next token involves sampling from this probability mass function to obtain a token $\hat{x}_{t+1}\sim p(x_{t+1}|\mathbf{x}_{t})$.
@@ 3.3 Differential Privacy
Definition 3.1 (Differential Privacy Dwork (2006)).
@@
The standard approach to integrating DP for deep learning models is by modifying the stochastic gradient descent algorithm (SGD), called DP-SGD Abadi et al. (2016). The modifications involve calculating per-sample gradients, rather than per-batch gradients, clipping them if their l-2 norm exceeds some threshold, then adding noise sampled from a Gaussian distribution. Note that per-sample gradient calculations typically reduce the parallel efficiency of GPUs; hence, applying DP-SGD has a negative impact on the training speed Yousefpour et al. (2021). In the following section, we will make use of DP-SGD.
@@ 4.1 Problem Formulation
Given a task-specific private dataset $D$, a pre-trained, generative teacher model $T$, a pre-trained, generative student model $S$, and a privacy budget $(\epsilon,\delta)$, our goal is to produce a student model such that $S$ minimizes Equation 3.1, the size of $S$ is significantly smaller than $T$, and $S$ is $(\epsilon,\delta)$-DP with respect to $D$. We will also assume access to a set of control codes $C$ Keskar et al. (2019) in our solution.
The DPKD framework introduced in Mireshghallah et al. (2022) tackles this problem setup by first fine-tuning a teacher model with DP-SGD, and then using DP-SGD a second time to minimize Eq 3.1 for the student. This solution has three limitations: (1) using DP-SGD on the student model results in additional error proportional to the total number of parameters of the student model, in the worst case Li et al. (2022); (2) allocating the privacy budget between the teacher and student can make it difficult for the student to obtain useful signal for strict privacy settings, smaller $\epsilon$ values; (3) larger runtime and memory consumption with an additional use of DP-SGD.
Rather than applying DP-SGD twice, we take the view that only applying DP-SGD on the teacher model leads to a more optimal solution. The framework from Yu et al. (2023a) follows this approach, but, as our results will show, the inclusion of more information about the teacher, such as the output distribution, substantially improves the utility of the student.
@@
8:Train Student $S=\min\limits_{S^{\prime}}\mathcal{L}_{\text{KD}}(T,S^{\prime},\tilde{D})$
@@ 4.2 DistilDP: Knowledge Distillation with DP Synthetic Text Data
We now introduce DistilDP, described in Algorithm 1. DistilDP fine-tunes a pre-trained, generative LLM using DP-SGD as our teacher model $T$, i.e., $T$ is $(\epsilon,\delta)$-DP. Afterward, $T$ is used to generate synthetic text data. In practice, we can use any synthetic text generation methodology, but we choose to follow an approach from Yue et al. (2022). In particular, we will utilize control codes Keskar et al. (2019) to provide more explicit control over text generation. Hence, sampling the next token $\hat{x}_{t+1}$ involves conditioning the model with a control code $c$ from a set of control codes $C$, in addition with the context vector $\mathbf{x}_{t}$: $\hat{x}_{t+1}\sim p(x_{t+1}|\mathbf{x}_{t},c)$.
@@ 4.3 Privacy Analysis
Proof. Our goal is for DistilDP to guarantee $(\epsilon,\delta)$-DP for $D$, which composes of text content for each data sample. After step 1, our method produces a teacher model $T$ that is $(\epsilon,\delta)$-DP. In step 2, sampling from $T$ incurs no extra privacy loss due to the post-processing property. However, one subtle, but crucial, detail is the potential privacy leakage from the control codes $C$. We limit the control codes to only contain categorical information of a data sample, not text contents. However, the categorical distribution in the original dataset may be private information. In our work, we follow the same procedure as Yue et al. (2022) by ignoring this low-cost privacy loss and using the exact categorical distribution. For step 3, since $\tilde{D}$ and $T$ are $(\epsilon,\delta)$-DP with respect to $D$, utilizing a non-private optimizer to minimize Equation 3.1 will result in $S$ being $(\epsilon,\delta)$-DP due to the post-processing property. ∎
@@ 5.1 Experimental Setup
We experimentally evaluated the privacy-utility tradeoff of our framework with the following setup. We used GPT2-Large Radford et al. (2019) and DistilGPT2 Sanh et al. (2019) as our teacher and student model, respectively. GPT2-Large contains 774M parameters while DistilGPT2 contains 82M parameters, so the teacher model is about 9.5 times larger than the student. Our implementation of fine-tuning language models with DP and knowledge distillation utilizes the dp-transformers library Inan et al. (2022), which combines the NLP library Huggingface Wolf et al. (2019) with the DP library Opacus Yousefpour et al. (2021).
We experimented with the Yelp Open dataset  [^1] https://www.yelp.com/dataset, which contains review text data on businesses. We follow the experimental setup as Yue et al. (2022) by using the review stars and business category as the control codes. These control codes are constructed as "Business Type: Restaurant |||| Review Stars: 3.0" and are prepended to each sample. The top 10 frequent business categories are sampled and reviews with no ratings are removed. The resulting data split is $1.9$M reviews for training, $5000$ for validation, and $5000$ for testing. Each data sample is set to a sequence length of $128$.
We also experimented with the Big Patent dataset Sharma et al. (2019), which consists of 1.3 million records of U.S. patent documents along with human-written abstractive summaries. For our experiments, we used the abstracts as the text data. To construct the control codes, we prepend each data sample with the Cooperative Patent Classification (CPC) code that it falls under. We only used a subset of the total number of documents, resulting in 150K abstracts for the training, 5000 for the validation, and 5000 for the testing dataset.
Lastly, we experimented on the DBpedia dataset Zhang et al. (2015), which is an ontology classification dataset consisting of 14 non-overlapping classes from DBpedia 2014. Each of these 14 ontology classes are used as a control code for the training samples. In total, we randomly selected 195K training samples, 5000 samples for validation, and 5000 samples for the test set.
We compare our method against three baselines: (1) a student model fine-tuned directly with DP-SGD, namely the teacher plays no role in training the student; (2) a student model fine-tuned using DPKD Mireshghallah et al. (2022); (3) a student model fine-tuned only on DP synthetic data (DP Syn Data), i.e. $\lambda=0$ in Eq. 3.1, which follows the framework of Yu et al. (2023a). The training hyperparameters used for our method and the baselines with the teacher and student models are shown in table 1. We set the privacy budget to $\epsilon=2$, which is a strong privacy guarantee. We follow the standard practice of setting $\delta=1/N$ where $N$ is the size of the dataset.
@@
For the generation of synthetic text, we use the truncation decoding strategies of top-$k$ Fan et al. (2018) and top-$p$ Holtzman et al. (2019) sampling using $k=50$ and $p=0.9$. The teacher model generates 400K synthetic samples using control codes from the Yelp and DBpedia training dataset and 200K synthetic samples from Big Patent. The maximum sequence length of each generated sequence is $128$. We follow the methodology of Kurakin et al. (2023) for training on DP synthetic data by splitting the Yelp and DBPedia synthetic samples into $396$K for the training set and $4$K for the validation. Likewise, $198$K and $2$K Big Patent synthetic examples for the training and validation set, respectively. For both our method and DPKD, we set $t=1.0$ and $\lambda=0.4$ for Equation 3.1. One minor technicality of our method is that because our teacher model was trained on control codes, its output distribution on the initial context vectors will be skewed toward the types of control codes. Hence, we ignore the first few softmax predictions in the distillation loss for each sample. We evaluated the utility of each method by measuring the perplexity (PPL) of the test set.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have accomplished amazing feats, from generating human-like text to improving human productivity through AI-powered assistants. However, recent work has demonstrated that practical privacy attacks are possible due to memorization of training data from LLMs carlini2019secret; carlini2021extracting. Differential Privacy (DP), a rigorous mathematical framework, is widely used to mitigate this type of information leakage dwork2006differential. To guarantee DP for machine learning models, DP is operationalized within the learning algorithm, known as DP-SGD abadi2016deep. The standard training paradigm of LLMs is to first pre-train an LLM on large publicly available data, then fine-tune it on a specific downstream task using private data. Recent works in differentially private training have taken advantage of this paradigm by initializing an LLM using pre-trained weights and then applying DP-SGD in the fine-tuning process, which can significantly boost its performance li2021large; yu2021differentially; ganesh2023public, albeit with some reduction in model utility.
Due to the huge computational and memory demands of LLMs, it becomes necessary to consider model compression in this training paradigm. For example, model compression is required when deploying LLMs on client devices where memory resources are scarce, or real-time systems such as sentence completion in text editors where inference latency is crucial xu2023federated. One popular technique for model compression that our work focuses on is task-specific Knowledge Distillation (KD) hinton2015distilling, which distills the knowledge of a larger teacher model onto a compressed student model. Although knowledge distillation in LLMs has been quite successful jiao2019tinybert; sanh2019distilbert; sun2019patient, few works have explored compressing LLMs with DP mireshghallah2022differentially; yu2023training; yu2023selective. Since achieving DP or KD for an LLM already trade-offs utility loss, accomplishing both can result in substantial utility loss, which highlights the challenge of this problem.
@@
A naive solution to this problem would be to apply DP-SGD when training the teacher model, then apply DP-SGD again during the knowledge distillation between the teacher and student. Since the teacher model is a function of the entire private corpus, applying DP-SGD during the fine-tuning of the teacher is seemingly unavoidable. Despite the additive computational and memory costs due to an additional run of DP-SGD, this solution has also been shown to have worst utility compared to differentially private fine-tuning the student without knowledge distillation mireshghallah2022differentially.
@@ 2 Related Works
Knowledge Distillation and DP Compression. Work in knowledge distillation for LLMs typically trains a smaller student model to match its output logits with the output logits of a larger, teacher model. More advanced works align the hidden state vectors between the teacher and the student jiao2019tinybert; sanh2019distilbert; sun2019patient.
@@
DP-SGD abadi2016deep is the most widely used technique to guarantee DP for LLMs. Specifically, a pre-trained LLM trained on large amounts of public text data scraped from the internet is further fine-tuned on a private corpus $D$ using DP-SGD mcmahan2017learning; yu2021differentially; li2021large. An orthogonal approach investigated differentially private decoding of LLMs flemings2024differentially; majmudar2022differentially; ginart2022submix, thereby avoiding the use of DP-SGD.
@@
mireshghallah2022differentially first introduced DP model compression and gave a framework for DP knowledge distillation. However, their framework requires using DP-SGD twice: once for the teacher and once for the student. This procedure leads to suboptimal results for text classification tasks. wang2019private also explored private model compression, but is limited by the number of queries to the teacher for knowledge distillation. yu2023selective explores DP model compression during the pre-training stage; our work focuses solely on compression during fine-tuning. Another line of work similar to our approach is Data-Free Distillation chen2019data; haroush2020knowledge where a generative model synthesizes data to substitute training data. Since no private data is used, this results in no privacy loss, $\epsilon=0$. Our approach uses a private corpus to generate synthetic data in a differentially private manner.
@@
Another problem setup broadly related to DP knowledge distillation is private ensemble learning via PATE papernot2016semi; papernot2018scalable. An ensemble of teacher models is trained on disjoint private data, then produces noisy labels on unlabeled public data, which comes from a similar distribution as the private data, for the student to train on. lyu2020differentially used PATE to achieve differentially private knowledge distillation for mobile analytics. Our method does not require unlabeled public data.
@@
DP Synthetic Data for Downstream Task. Promising progress has been made in generating high-fidelity, DP synthetic text bommasani2019towards; mattern2022differentially; kurakin2023harnessing; yue2022synthetic. These works have substantially reduced the utility gap between DP-SGD on real data and SGD on DP synthetic data for classification tasks.
@@
yu2023training investigated synthetic data generation from DP foundational models to train lightweight downstream models. However, our method differs from theirs by the following: (1) Their framework only considers knowledge transfer from the synthetic data. Our results demonstrate that other forms of knowledge distillation, in particular, the output distribution and hidden representations, are crucial to obtain strong performance for compressed models. (2) Their framework needs large foundational models to generate high-enough quality synthetic text data, which requires model parallelism on four V100 GPUs for fine-tuning even with LoRA hu2021lora; yu2021differentially. And to the best of our knowledge, there is no open-source implementation of DP model parallel training. In contrast, our framework can obtain strong results without relying on large foundational models and thus does not need model parallelism.
@@ 3.1 Knowledge Distillation
Knowledge Distillation (KD) aims at transferring the knowledge of a teacher LLM $T$ to a smaller student LLM $S$. The standard approach is to train the student to mimic the output distribution of the teacher and the true labels. Let $f_{T}(x)$ and $f_{S}(x)$ be the logit outputs of the teacher and student networks evaluated on an input data $x$, respectively, and $\mathbf{y}$ be the true labels. The student is trained to minimize the linear combination of the supervised learning loss and the distillation loss:
@@
where $\lambda$ is a weight hyperparameter, $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, $\mathcal{L}_{\text{KL}}$ is the Kullback-Leibler divergence, $\sigma$ is the softmax function, and the second term is the distillation loss which follows the softmax-temperature using a temperature parameter $t$ hinton2015distilling. At inference time, $t$ is set to 1 to recover the standard softmax. y is known as the hard label because of its one-hot encoding while $\sigma(f_{T}(x)/t)$ is the soft label.
@@ 3.2 Text Generation
Given some context vector $\mathbf{x}_{t}=x_{1},x_{2},...,x_{t}$, which is a string of tokens from some vocabulary $V$, i.e. $x_{i}\in V$ for all i=1,,ti=1,...,t, text generation involves sampling the next token $x_{t+1}$ using a generative language model $p$. More precisely, the output of a language model $p$ for a given context $\mathbf{x}_{t}$ is a likelihood function of all possible tokens $p(x_{t+1}=w|\mathbf{x}_{t})$, and choosing the next token involves sampling from this probability mass function to obtain a token $\hat{x}_{t+1}\sim p(x_{t+1}|\mathbf{x}_{t})$.
@@ 3.3 Differential Privacy
Definition 3.1 (Differential Privacy dwork2006differential).
@@
The standard approach to integrating DP for deep learning models is by modifying the stochastic gradient descent algorithm (SGD), called DP-SGD abadi2016deep. The modifications involve calculating per-sample gradients, rather than per-batch gradients, clipping them if their l-2 norm exceeds some threshold, then adding noise sampled from a Gaussian distribution. Note that per-sample gradient calculations typically reduce the parallel efficiency of GPUs; hence, applying DP-SGD has a negative impact on the training speed yousefpour2021opacus. In the following section, we will make use of DP-SGD.
@@ 4.1 Problem Formulation
Given a task-specific private dataset $D$, a pre-trained, generative teacher model $T$, a pre-trained, generative student model $S$, and a privacy budget $(\epsilon,\delta)$, our goal is to produce a student model such that $S$ minimizes Equation 3.1, the size of $S$ is significantly smaller than $T$, and $S$ is $(\epsilon,\delta)$-DP with respect to $D$. We will also assume access to a set of control codes $C$ keskar2019ctrl in our solution.
The DPKD framework introduced in mireshghallah2022differentially tackles this problem setup by first fine-tuning a teacher model with DP-SGD, and then using DP-SGD a second time to minimize Eq 3.1 for the student. This solution has three limitations: (1) using DP-SGD on the student model results in additional error proportional to the total number of parameters of the student model, in the worst case li2022does; (2) allocating the privacy budget between the teacher and student can make it difficult for the student to obtain useful signal for strict privacy settings, smaller $\epsilon$ values; (3) larger runtime and memory consumption with an additional use of DP-SGD.
Rather than applying DP-SGD twice, we take the view that only applying DP-SGD on the teacher model leads to a more optimal solution. The framework from yu2023training follows this approach, but, as our results will show, the inclusion of more information about the teacher, such as the output distribution, substantially improves the utility of the student.
@@
[…]
@@ 4.2 DistilDP: Knowledge Distillation with DP Synthetic Text Data
We now introduce DistilDP, described in Algorithm 1. DistilDP fine-tunes a pre-trained, generative LLM using DP-SGD as our teacher model $T$, i.e., $T$ is $(\epsilon,\delta)$-DP. Afterward, $T$ is used to generate synthetic text data. In practice, we can use any synthetic text generation methodology, but we choose to follow an approach from yue2022synthetic. In particular, we will utilize control codes keskar2019ctrl to provide more explicit control over text generation. Hence, sampling the next token $\hat{x}_{t+1}$ involves conditioning the model with a control code $c$ from a set of control codes $C$, in addition with the context vector $\mathbf{x}_{t}$: $\hat{x}_{t+1}\sim p(x_{t+1}|\mathbf{x}_{t},c)$.
@@ 4.3 Privacy Analysis
Proof. Our goal is for DistilDP to guarantee $(\epsilon,\delta)$-DP for $D$, which composes of text content for each data sample. After step 1, our method produces a teacher model $T$ that is $(\epsilon,\delta)$-DP. In step 2, sampling from $T$ incurs no extra privacy loss due to the post-processing property. However, one subtle, but crucial, detail is the potential privacy leakage from the control codes $C$. We limit the control codes to only contain categorical information of a data sample, not text contents. However, the categorical distribution in the original dataset may be private information. In our work, we follow the same procedure as yue2022synthetic by ignoring this low-cost privacy loss and using the exact categorical distribution. For step 3, since $\tilde{D}$ and $T$ are $(\epsilon,\delta)$-DP with respect to $D$, utilizing a non-private optimizer to minimize Equation 3.1 will result in $S$ being $(\epsilon,\delta)$-DP due to the post-processing property. ∎
@@ 5.1 Experimental Setup
We experimentally evaluated the privacy-utility tradeoff of our framework with the following setup. We used GPT2-Large radford2019language and DistilGPT2 sanh2019distilbert as our teacher and student model, respectively. GPT2-Large contains 774M parameters while DistilGPT2 contains 82M parameters, so the teacher model is about 9.5 times larger than the student. Our implementation of fine-tuning language models with DP and knowledge distillation utilizes the dp-transformers library inan2022dp-transformers, which combines the NLP library Huggingface wolf2019huggingface with the DP library Opacus yousefpour2021opacus.
We experimented with the Yelp Open dataset  [^1] https://www.yelp.com/dataset, which contains review text data on businesses. We follow the experimental setup as yue2022synthetic by using the review stars and business category as the control codes. These control codes are constructed as "Business Type: Restaurant $|$ Review Stars: 3.0" and are prepended to each sample. The top 10 frequent business categories are sampled and reviews with no ratings are removed. The resulting data split is 1.91.9M reviews for training, $5000$ for validation, and $5000$ for testing. Each data sample is set to a sequence length of $128$.
We also experimented with the Big Patent dataset sharma2019bigpatent, which consists of 1.3 million records of U.S. patent documents along with human-written abstractive summaries. For our experiments, we used the abstracts as the text data. To construct the control codes, we prepend each data sample with the Cooperative Patent Classification (CPC) code that it falls under. We only used a subset of the total number of documents, resulting in 150K abstracts for the training, 5000 for the validation, and 5000 for the testing dataset.
Lastly, we experimented on the DBpedia dataset zhang2015character, which is an ontology classification dataset consisting of 14 non-overlapping classes from DBpedia 2014. Each of these 14 ontology classes are used as a control code for the training samples. In total, we randomly selected 195K training samples, 5000 samples for validation, and 5000 samples for the test set.
We compare our method against three baselines: (1) a student model fine-tuned directly with DP-SGD, namely the teacher plays no role in training the student; (2) a student model fine-tuned using DPKD mireshghallah2022differentially; (3) a student model fine-tuned only on DP synthetic data (DP Syn Data), i.e. $\lambda=0$ in Eq. 3.1, which follows the framework of yu2023training. The training hyperparameters used for our method and the baselines with the teacher and student models are shown in table 1. We set the privacy budget to $\epsilon=2$, which is a strong privacy guarantee. We follow the standard practice of setting $\delta=1/N$ where $N$ is the size of the dataset.
@@
For the generation of synthetic text, we use the truncation decoding strategies of top-$k$ fan2018hierarchical and top-$p$ holtzman2019curious sampling using $k=50$ and $p=0.9$. The teacher model generates 400K synthetic samples using control codes from the Yelp and DBpedia training dataset and 200K synthetic samples from Big Patent. The maximum sequence length of each generated sequence is $128$. We follow the methodology of kurakin2023harnessing for training on DP synthetic data by splitting the Yelp and DBPedia synthetic samples into 396396K for the training set and 44K for the validation. Likewise, 198198K and 22K Big Patent synthetic examples for the training and validation set, respectively. For both our method and DPKD, we set $t=1.0$ and $\lambda=0.4$ for Equation 3.1. One minor technicality of our method is that because our teacher model was trained on control codes, its output distribution on the initial context vectors will be skewed toward the types of control codes. Hence, we ignore the first few softmax predictions in the distillation loss for each sample. We evaluated the utility of each method by measuring the perplexity (PPL) of the test set.